Hi, please help...
A family has two cars. In a given week, the first car gets an average of 30 miles per gallon, and the second car gets 25 miles per gallon. The two cars combined drive a total of 1525 miles in that week, for a total gas consumption of 55 gallons. How many gallons were consumed by each of the two cars that week?

Answer by Alan3354(69443)

A= distance driven by 1st car (30 mpg)
1st car used A/30 gallons
B= distance driven by 2nd car (25 mpg)
2nd car used B/25 gallons
A+B = 1525
A/30 + B/25 = 55
A = 1525-B
(1525-B)/30 + B/25 = 55
Multiply by 150 to clear fractions
5(1525-B) + 6B = 8250
7625-5B + 6B = 8250
B + 7625 = 8250
B = 625 miles
A = 900 miles